Two parallel wires are separated by a distance of 0.05 m. Wire 1 carries a


carries current of 12.0 A. Wire 2 carries a current of 6.0 A flowing in the


opposite direction as shown in the illustration.


Find the


(a) magnitude and


direction of the magnetic field created by wire 1 at wire 2,



(b) force exerted


by wire 1 on the 2.0 m long section of wire 2, and



(c) force exerted by wire 2


on wire 1.

1
Expert's answer
2022-04-11T12:04:51-0400

a)

"B_1=\frac{\mu_0I_1}{2\pi d}=48~\mu T,"

b)

"F_{12}=\frac{\mu_0I_1I_2 l}{2\pi d}=0.58~mA,"

c)

"F_{21}=\frac{\mu_0I_2I_1 l}{2\pi d}=0.58~mA."